Toyota Motor Corp. Us 2017 annual report

How does the company's gross profit percentage compare to the industry? What are the company's inventory turnover and days sales in inventory for the year end? Is there anything interesting to note about these ratios?
Answer #1

toyota gross profit 2017=gross profit/sales
=4862857000/27597193000
=17.62%
The industry gross profit is 10% to 15% and toyota is doing better than the industry average.
Inventory turnvoer=cost of goods sold/inventory
=22734336000/2388617000=9.52
Days sales inventory=365/inventory turnvoer
=365/9.52=38.35 days
The inventory it is holding is less than the industry average since it follows many methods which reduces its days sales inventory